OpenStudy (anonymous):

a stone is dropped from a captive balloon at an elevation of 1000ft,. two sec later another stone is projected vertically upward from the ground with a velocity of 248ft per sec. if g is 32ft per sec squared when and where will the stones pass each other?

razor99 (razor99):

wow is that a physics question

OpenStudy (anonymous):

obviously..yes :)

razor99 (razor99):

this is the first time i ever saw a physics question lol cuz i study business stream

OpenStudy (anonymous):

can u write a equation for the first stone distance travelled by it in t seconds

OpenStudy (anonymous):

i know that: \[V _{01}=0 ft/s\] \[V _{02}=248 ft/s\] \[g=32.2 ft/s^2\] \[t _{1}=t _{2}+2\]

OpenStudy (anonymous):

@unkabogable can u?

OpenStudy (anonymous):

So for the first stone distance travelled by it in t seconds is = ut + 1/2 at^2 = 16t^2

OpenStudy (anonymous):

right?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

now find, for the first stone distance travelled by it in 2 seconds

OpenStudy (anonymous):

can u?

OpenStudy (anonymous):

yeah...wait: so i have these eq. \[Y _{1}=1000-Y _{2}\] \[Y _{2}=V _{02}t _{2}-1/2(32.2)t _{2}^2\] then i'll substitute and i'll verify my answer later

OpenStudy (anonymous):

Let me do it then, for the first stone distance travelled by it in 2 seconds=16*2^2 = 64ft now velocity of it after two seconds = (o^2 + 2*32 *64)^1/2 = 64ft/s

OpenStudy (anonymous):

Now,Now, after two seconds, height of the stone after t seonds =1000-64-( ut+1/2 at^2) = 936-64t - 16t^2

OpenStudy (anonymous):

then for the next stone: height of the stone after t seconds = ut + 1/2 at^2 = 248t + 1/2 (-32) t^2 =248t - 16t^2

OpenStudy (anonymous):

Now, when the the stones pass each other , 936-64t - 16t^2=248t - 16t^2 936=312t t=3seconds

OpenStudy (anonymous):

so the stone will pass each other at 3+2 seconds = 5 seconds

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes..thanks again!! :)

OpenStudy (anonymous):

WElcome

OpenStudy (anonymous):

No, can u find this where will the stones pass each other

OpenStudy (anonymous):

Now

OpenStudy (anonymous):

(248)(5-2)\[(248)(5-2)- \frac{ (32.2(5-2)^2) }{ 2} = 600 ft\] ??

OpenStudy (anonymous):

Yep........ 600 ft above the ground